OpenStudy (anonymous):

Any1 online here that can help me out?

OpenStudy (anonymous):

Yes go on

OpenStudy (anonymous):

What is the final concentration of (in M) 16.6 mL iron(III) chloride solution with an initial concentration of 1.343 M which is diluted with 15.2 mL of solvent? Assume the volumes are additive.

OpenStudy (anonymous):

no of milli moles of solute = (16.6)(1.343) On dilution, no of milli moles of solute remain constant, the volume of solution changes new volume of solution = 16.6 mL+ 15.2 mL = 31.8 mL Therefore, final concentration = no of milli moles of solute / new volume of solution = 16.6*1.343 / 31.8 = 0.701 M

OpenStudy (anonymous):

I freaking love you :) Thank you so much. Actually makes sense now. Also, if i get stuck again, mind if i ask you for help again?

OpenStudy (anonymous):

No problem :)